given logx+y/log2 = log x-y/log3 = log64/log 0.125
find the values of x and y


please do reply of d above ques...

Answer by Alan3354(21583)

logx+y/log2 = log x-y/log3 = log64/log 0.125
----------
logx+y/log2 = log64/log 0.125
Cross multiply
log(x+y)*log(0.125) = log2*log64
log(x+y) = log2*log64/log(0.125)
log(x+y) =~ -0.60206
x+y = 0.25
-------------------------------
log x-y/log3 = log64/log 0.125
Cross multiply, etc
log(x-y) = log3*log64/log(0.125)
log(x-y) =~
x-y =~ 0.95424
x-y = 1/9
x+y = 1/4
----------- Add
2x = 13/36
x = 13/72
---------
y = 5/72